OpenStudy (anonymous):

A pie is removed from a 375°F oven and cools to 215°F. after 15 minutes in a room at 72°F. How long (from the time it is removed from the oven) will it take the pie to cool to 120 f ?

OpenStudy (kira_yamato):

By Newton's Law of Cooling: dT/dt = k(T - 72) -160 = 303k k = -160/303 dt/dT = 1/k * 1/(T - 72) t + c = 1/k * ln |T - 72| k(t+c) = ln|T - 72| T = 72 + Ae^kt, where A = ±e^c When t = 0, T = 375, A = 303 T = 303 e^kt When T = 120, kt = ln (120/303) t = -(160ln (120/303))/303

OpenStudy (anonymous):

ty